# Near-Inertial Pollard Waves Modeling the Arctic Halocline

### v3

@@ 2.3. North Pole and coordinates relative to the TDC
with basis given by $\mathbf{e}_{\varphi}^{{\dagger}}(\mathfrak{N}),\mathbf{e}_{\theta}^{{\dagger}}((\mathfrak{N}))$, with $\mathfrak{N}$ representing the North Pole. See Figure 2.2 for a depiction of the basis at the North Pole.
[…]
Figure $2.$ Depiction of the basis at the North Pole. In reality, the angle $\gamma$ is bigger than the one depicted here (close to $90^{\circ}$). However, for graphical clarity reasons, we did it smaller ($70^{\circ}).$
@@
These unit vectors are given by (see Figure 2.2)
@@ 3. Description of the flows
As depth increases, turbulent effects generated by momentum exchange at the water surface rapidly decay. This attenuation is even more pronounced in the presence of sea ice, a defining feature of the oceanic region considered here. For a detailed analysis of wind–ice–water interactions at the Arctic Ocean surface, we refer to [MIO] where a solution for nonlinear wind-induced ice-drift currents is found, featuring the superposition of near-inertial oscillations, a mean Ekman flow and a geostrophic current [^1] We remark that, for studies of ice–sea or air–sea interactions, the appropriate governing equations are the Navier–Stokes equations rather than the Euler equations, which are used in this work, as we focus on the deeper parts of the ocean..
@@
which prevents the exchange of fluid particles between adjacent layers (see [ConstantinBook]). This condition reflects the strong stratification of the halocline, which inhibits direct contact between the warmer, saltier AW and the sea ice above.

### v4

@@ 2.3. North Pole and coordinates relative to the TDC
with basis given by $\mathbf{e}_{\varphi}^{{\dagger}}(\mathfrak{N}),\mathbf{e}_{\theta}^{{\dagger}}((\mathfrak{N}))$, with $\mathfrak{N}$ representing the North Pole. See Figure 4 for a depiction of the basis at the North Pole.
Figure 4. Depiction of the basis at the North Pole. In reality, the angle $\gamma$ is bigger than the one depicted here (close to $90^{\circ}$). However, for graphical clarity reasons, we did it smaller ($70^{\circ}).$
@@
These unit vectors are given by (see Figure 4)
@@ 3. Description of the flows
As depth increases, turbulent effects generated by momentum exchange at the water surface rapidly decay. This attenuation is even more pronounced in the presence of sea ice, a defining feature of the oceanic region considered here. For a detailed analysis of wind–ice–water interactions at the Arctic Ocean surface, we refer to [50] where a solution for nonlinear wind-induced ice-drift currents is found, featuring the superposition of near-inertial oscillations, a mean Ekman flow and a geostrophic current [^1] We remark that, for studies of ice–sea or air–sea interactions, the appropriate governing equations are the Navier–Stokes equations rather than the Euler equations, which are used in this work, as we focus on the deeper parts of the ocean..
@@
which prevents the exchange of fluid particles between adjacent layers (see [8]). This condition reflects the strong stratification of the halocline, which inhibits direct contact between the warmer, saltier AW and the sea ice above.